OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the syste

OpenStudy (anonymous):

please help me

OpenStudy (marissalovescats):

Hey there

OpenStudy (anonymous):

hi

OpenStudy (marissalovescats):

Well since we are solving for x, we need to eliminate the ys with elimination.

OpenStudy (anonymous):

yea i need help on that

OpenStudy (anonymous):

I think that x = 1

OpenStudy (marissalovescats):

What can we multiply the top equation with to cancel out the +2y in the bottom equation?

OpenStudy (anonymous):

okay did it

OpenStudy (marissalovescats):

So what did you multiply the top equation by?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

1)2(3x + y = 6) 2) 6x + 2y = 8 , 6x+2y=12 - 6x + 2y = 8 ___________________ 0 =4 , so no solution

OpenStudy (marissalovescats):

Okay multiplying the top equation by 6 gives us this  Hm... Yes @ASAAD123 is correct because any term that you multiply the top and bottom with to cancel out the y, it also cancels the x. Therefore you cannot solve for x. X has no solution. I was getting there :P